OpenStudy (anonymous):

f’’’-f’^2+1=0 solution of this equation with boundary conditions f’(0)=0, f’(∞)=1, f’’(∞)=0 is needed so badly

OpenStudy (alekos):

have you attempted this?

OpenStudy (anonymous):

no

OpenStudy (kc_kennylau):

1. which grade are you in?

OpenStudy (kc_kennylau):

2. f'^2 means \([f'(x)]^2\) or \(f''(x)\)?

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

square of f'

OpenStudy (alekos):

Is the first term f''' or f''

OpenStudy (kc_kennylau):

This is too complicated, this should not be what you'll be asked... can you double check your question?

OpenStudy (anonymous):

question double checked :) the first one is f'''

OpenStudy (kc_kennylau):

still the third degree?

OpenStudy (alekos):

i agree, this is not a standard question

OpenStudy (anonymous):

yes

OpenStudy (kc_kennylau):

Please consult your teacher.

OpenStudy (anonymous):

this is my homework for phd lecture advanced math

OpenStudy (anonymous):

i know it s complicated that s why searching for help

OpenStudy (alekos):

can you give me a day or two to come up with a solution?

OpenStudy (anonymous):

of course that would be very nice

OpenStudy (anonymous):

due is two days from now

OpenStudy (anonymous):

If you find a solution please post it here too. Third order non-linear DE... I think you need to be a real maths wiz to solve this.

OpenStudy (kagıtucak):

you need to solve this as differential equation from the charactreristic equation r^3-2r+1=0 and clearly r=1 is a root for this then (r-1)(r^2+r-1)=0 other roots will come from r^2+r-1=0 (r+1/2)^2-5/4=0 (r+1/2)=square root of 5/4 r1=-1/2+ (root5)/2 and r2=-1/2-(root 5)/2 then your solution must be in the form of y(t)=a.e^t+b.e^r1+c.e^r2 then take its first derivative and it will equal to 0 at t=0 and 1 at t=infinity from the given information

OpenStudy (anonymous):

thanks i ll give it a shot

OpenStudy (amoodarya):

do you solve it by computer ? or handy ? if you must solve it handy you can use variational method like( ptero - petro galarkin method ) or if not you want approximate solution suppose f(x)=a+bx+cx^2+dx^3 put them in equation and find a,b,c,d any way another method is solving by series (because it is not a linear ode )

OpenStudy (anonymous):

i ve to solve it by hand, will consider recommendations

OpenStudy (amoodarya):

are you in bs level or ms level "sinan"?

OpenStudy (anonymous):

ms level

OpenStudy (kagıtucak):

amoodarya is wrong f(x) cannot be in the form of as he said because when we take its derivative and put infinity we get infinity not 1 so it is absolutely wrong

OpenStudy (amoodarya):

ok so boundary conditions f’(∞)=1, f’’(∞)=0 are not usual boundary conditions in ms level , because you have not learn ode course !!! any way by this boundary conditions f’(∞)=1, f’’(∞)=0 you need a function like exp(-x), exp(-2x) ,... so in ∞ goes to zero .

OpenStudy (kagıtucak):

exactly

OpenStudy (anonymous):

r^3-2r+1=0 is the characteristic equation for f'''-f'^2+1=0 ? ?

OpenStudy (amoodarya):

no it is not the characteristic equation for f'''-f'^2+1=0

OpenStudy (anonymous):

because it is square of f'

OpenStudy (anonymous):

i couldn t even write characteristic equation

OpenStudy (amoodarya):

r^3-2r+1=0 is for f'''=2f'+f=0

OpenStudy (anonymous):

yes

OpenStudy (kagıtucak):

you are right i didnt realize that

OpenStudy (amoodarya):

i got an idea !

OpenStudy (anonymous):

seems like a good idea, what to do after ?

OpenStudy (amoodarya):

i have been thinking about the integration of g'

OpenStudy (anonymous):

Did you try laplace transforms? lol

OpenStudy (alekos):

Looking at laplace there is no easy way to do a transform for f'^2 Looking at homogenous ODE's there is no characteristic eqn for f'^2 I cant see a way forward to a solution

OpenStudy (anonymous):

this question is a real challange :)

OpenStudy (anonymous):

the general solution is \[f(x)=Ae ^{x}+Be ^{\frac{ -1+\sqrt{5} }{ 2 }x}+Ce ^{\frac{ -1-\sqrt{5} }{ 2 }x}\] and use the boundary condition to find A,B,C

OpenStudy (anonymous):

thanx but how did you come to this ?

OpenStudy (anonymous):

from charactreristic equation \[m ^{3}-2m+1=0 \] \[(m-1)(m ^{2}+m-1)=0\]

OpenStudy (kc_kennylau):

it's not 2m. it's f'^2 not 2f'

OpenStudy (anonymous):

yes kc is true

OpenStudy (anonymous):

thats the hard part f'^2

OpenStudy (anonymous):

ah srry

OpenStudy (alekos):

if we let g=f' we get g''-g^2+1=0 according to my research this DE solves to an elliptic integral In general, elliptic integrals cannot be expressed in terms of elementary functions, so there may be no solution to this particular problem

OpenStudy (anonymous):

I was thinking of saying f'=p like yours then the equation becomes p.p''-p^2+1=0 after that if we say p''=z equation will be pz-p^2+1=0 and z= (p^2-1)/p

OpenStudy (anonymous):

do you think i can reach somewhere from here ?

OpenStudy (alekos):

if you make p=f' you'll get p''-p^2+1=0 which is essentially the same, so I don't think it'll work

OpenStudy (anonymous):

what if after getting z=(p^2-1)/p integrating both side

OpenStudy (kainui):

I was playing around noticing that if you plug in some values, you can find out from: f’(0)=0, f’(∞)=1, f’’(∞)=0 f’’’(∞)-f’(∞)^2+1=0 that: f’’’(∞)-(1)^2+1=0 f’’’(∞)=0 And you can take the derivative of the original formula d/dx(f’’’-f’^2+1=0) f’’’’-2f’*f’’=0 f’’’’=2f’*f’’ f’’’’(∞)=2f’(∞)*f’’(∞) f’’’’(∞)=0 and that will continue to lead us down the road that: \[f^{(n)}(∞) =0\] for n>1. Continuing using our differential equation we can see: f’’’(0)-f’(0)^2+1=0 f’’’(0)=-1 but the trail stops here since we don't know f’’(0). So from this it seems pretty fair and quite simple to guess from that: \[f’(x)=1-e^{-x}\] since it fulfills all our boundary conditions: f’(0)=0, f’’’(0)=-1; f’(∞)=1, f^(n)(∞)=0 for n>1 So then, fairly satisfied, we can then integrate this to get an answer that fulfills the boundary conditions. \[f(x)=x+e^{-x}+C\]

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

all yours :)

OpenStudy (anonymous):

note this is a first-order ODE in \(u=f'\):$$u'+u^2+1=0$$hence $$\frac{du}{dt}=-(1+u^2)\\\frac1{1+u^2}\frac{du}{dt}=-1\\\int \frac1{1+u^2}\frac{du}{dt}dt=\int-1\ dt$$by the chain-rule the left simplifies$$\int\frac1{1+u^2}du=\int -1\ dt\\\arctan u=c_1-t\\u=\tan(c_1-t)$$

OpenStudy (anonymous):

to solve for \(f\) we have that \(u=\frac{df}{dt}\) so$$\frac{df}{dt}=\tan(c_1-t)\\f(t)=\int \tan(c_1 -t)\ dt=-\log|c_1-t|+c_2$$

OpenStudy (anonymous):

actually I may have misread your initial equation -- but I'm too far in to stop so: f’’’-f’^2+1=0 solution of this equation with boundary conditions f’(0)=0, f’(∞)=1, f’’(∞)=0 is needed so badly $$u(t)=\tan(c_1-t)\\u(0)=0\implies c_1=\pi n,n\in\mathbb{Z}\\\lim_{t\to\infty}u(t)=1\ \text{oops! u(t) permits no such limit}$$

OpenStudy (alekos):

Nice try Oldrin. Kainui, if you substitute f'= 1-e^-x, f'' = e^-x, f''' = -e^-x in the original equation then you don't get zero on the LHS, so I don't think that your valiant attempt at a solution is correct.

OpenStudy (kainui):

@alekos haha yeah that's kind of a problem isn't it?

OpenStudy (alekos):

Sinan, I'd be really interested to know if your lecturer has a solution to this very perplexing problem

OpenStudy (kainui):

http://www.wolframalpha.com/input/?i=y''%3D(y%2B1)(y-1)&t=crmtb01 It says there's a step-by-step solution for this on Wolfram actually.

OpenStudy (anonymous):

something tells me elliptic functions are well past intro ODE material

OpenStudy (alekos):

Kainui, Sinan I just plugged in y'''-y'^2+1=0 in the Wolfram Alpha DE solver and we get a Weierstrass Zeta function which is an elliptic function. I've never come across this type of function before, and it does say that the step by step solution is available after creating an account.

OpenStudy (alekos):

Just created an account but it says that there is no step by step solution available? So we are none the wiser as to how this was arrived at, but at least we know the answer.

OpenStudy (anonymous):

yes i know i tried everything it s not a simple question, anyway i submitted whatever i could do :) i ll ask my lecturer if he has solution, i ll keep you posted :) thanx anyway